# LLMs Struggle to Perform Counterfactual Reasoning with Parametric Knowledge

### v1

@@ 5 Discussion
Our work highlights LLM challenges in counterfactual reasoning requiring dynamic integration of parametric and contextual knowledge. Experiments show LLMs often default to parametric knowledge or fail to balance contextual override with selective retrieval. Simple finetuning is limited, inducing shortcuts or degrading knowledge. Pretraining with counterfactual data, while improving such reasoning, can also harm factual task performance. These findings point to a core limitation: current LLMs lack robust mechanisms for on-the-fly, conditional use of their parametric knowledge.
6 Limitations
While our study sheds light on the challenges LLMs face in integrating parametric knowledge with novel counterfactual premises, it is subject to several limitations. Our synthetic benchmarks abstract away many complexities of real-world reasoning. In the toy setting, counterfactual premises are expressed as single-edge edits to a static knowledge graph and queries are limited to two-hop chains. Many real-world scenarios often involve multi-predicate interactions, ambiguous or probabilistic relationships, and noisy or conflicting evidence from multiple sources. What is significant is that we show how models struggle even under these simpler circumstances.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Full Prompt Library
‘<Cause>‘: The event assumed to be the cause in the causal-query prompt (e.g., “Rain”).
‘<Effect>‘: The event assumed to be the effect in the causal-query prompt (e.g., “Flooding”).
@@
A.2 A. General Causal-Query Prompts
@@ 1. Standard (Direct) Prompt
2. Chain-of-Thought (CoT) Prompt
@@
> **Fine-Tuned** model runs used these exact templates but pointed at the fine-tuned checkpoint.
@@ 2. Anticausal Chain Expansion
4. Post-Transition Chain Expansion
@@ Appendix E LLM Usage
Appendix F Compute

### v2

@@ 5 Discussion
Our work highlights LLM challenges in counterfactual reasoning requiring dynamic integration of parametric and contextual knowledge. Experiments show LLMs often default to parametric knowledge or fail to balance contextual override with selective retrieval. Simple finetuning is limited, inducing shortcuts or degrading knowledge. Pretraining with counterfactual data, while improving such reasoning, can also harm factual task performance. These findings point to a core limitation: current LLMs lack robust mechanisms for on-the-fly, conditional use of their internal knowledge. In practice, we observe two recurring failure modes—context-ignoring and context-overfitting—that persist across prompting strategies and fine-tuning. While our study sheds light on the challenges LLMs face in integrating parametric knowledge with novel counterfactual premises, it is subject to several limitations. In the toy setting, counterfactual premises are expressed as single‐edge edits to a static knowledge graph and queries are limited to two‐hop chains. Many real-world scenarios often involve multi‐predicate interactions, ambiguous or probabilistic relationships, and noisy or conflicting evidence from multiple sources. What is significant is that we show how models struggle even under these simpler circumstances. Future work could focus on extending our analysis to deeper and more noisy multi-hop relations.
@@ A.1 Full Prompt Library
‘<Cause>‘: The event assumed to be the cause in the causal‐query prompt (e.g., “Rain”).
‘<Effect>‘: The event assumed to be the effect in the causal‐query prompt (e.g., “Flooding”).
@@
A.2 A. General Causal‐Query Prompts
@@ 1. Standard (Direct) Prompt
2. Chain‐of‐Thought (CoT) Prompt
@@
> **Fine‐Tuned** model runs used these exact templates but pointed at the fine‐tuned checkpoint.
@@ 2. Anticausal Chain Expansion
4. Post‐Transition Chain Expansion
@@ Appendix E LLM Usage
Appendix F Toy Experiment Hyperparameters
| Hyperparameter | Values |
| Learning Rate | |
| Weight Decay | |
| Batch Size | 512 |
Appendix G Compute